OpenStudy (anonymous):

At a certain rate of simple interest, a certain sum doubles itself in 10 years. In how many years will it triple itself?

OpenStudy (lgbasallote):

15 years?

OpenStudy (anonymous):

How??

OpenStudy (anonymous):

I think your answer is correct lgbasallote, because 15 is one of the options.

OpenStudy (lgbasallote):

what i did was it doubles in 10 years...so it quadruples in 20...in between is 150 -> triple

OpenStudy (lgbasallote):

15*

OpenStudy (anonymous):

I there no other way?

OpenStudy (anonymous):

*Is

OpenStudy (lgbasallote):

i dont know...seems there is an algebraic equation for it..but i cant equate it..

OpenStudy (anonymous):

Can we take the rate as a constant(say, 5%) and the principal as x?

OpenStudy (lgbasallote):

ratio and proportion can be used too i guess 2:10 = 3:x 30 = 2x x = 15

OpenStudy (anonymous):

@amistre64 can you help?

OpenStudy (experimentx):

ln(2) = 10 ln(1+ r/100) => rate must be 7.17735 ln(3) = x ln(1+ 7.17735/100) => x = 15.8496 http://www.wolframalpha.com/input/?i=solve+ln%283%29+%3D+x+ln%281%2B+7.17735%2F100%29

OpenStudy (anonymous):

But its not compound interest, its simple interest.

OpenStudy (anonymous):

What is In?

OpenStudy (experimentx):

I = PTR/100 => I = P => R => 100/10 for tripling 2I = PTR/100 => T=15

OpenStudy (anonymous):

How is I = P(in the second step)??

OpenStudy (experimentx):

for sum to double .. . you need Interest = Principal => I+P = A = 2P and same logic goes for tripling

OpenStudy (anonymous):

Oh, I see... thanks.

OpenStudy (anonymous):

Sorry to trouble you but how is R = 100/10.

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

I and R cancel out I = PTR/100 =>1 = TR/100 => T = 10 => 1 = R10/100 => R = 100/10

OpenStudy (anonymous):

Thanks.